OpenStudy (anonymous):

Find all solutions to the equation sin(3x)cosx - sinx cos(3x) = 0 on the interval [0,2pi]. Check all that apply.

OpenStudy (anonymous):

a. x=0,pi,2pi b. x=0, pi/2 ,pi, 3pi/2, 2pi c. x= 0, pi/2, 3pi/2 d. x= 0, pi/4, pi/2, 3pi/4, pi, 2pi/2, 2pi

OpenStudy (anonymous):

I think it is B

OpenStudy (anonymous):

it can be 0=sin(3x-x); x=0||x=pi||x=2pi

OpenStudy (anonymous):

because sinacosb-sinbcosa=sin(a-b);

OpenStudy (anonymous):

It was B, but thank you.

OpenStudy (anonymous):

are you sure?